Math 3150-001: Partial Differential Equations for Engineers Spring 2012

Instructor: Andrej Cherkaev, JWB 225, 581-6822, cherk@math.utah.edu.

Time and Place: M., W., 10:45-11:35, WEB L102

Office Hours: F 10:00-12:00 pm, or by appoitment.

Prerequisite: Calculus series, ODE, Linear Algebra

Text: Partial Differential Equations with Fourier Series and Boundary Value Problems by Nakhle Asmar, Prentice-Hall, 2005; Second Edition

Sections: 1.1, 1.2; 2.1 - 2.4; 3.1 - 2.9; 4.1 - 4.4; 7.1 - 7.4

1  1.1-1.2  Introduction
2  2.1-2.2  Periodic Functions and Fourier Series 3
2.3-2.4  More on Fourier Series 4  3.1
Examples in Physics and Engineering, Midterm 1 5  3.2-3.3  One Dimensional Wave Equation 6
3.4  DAlemberts Method 7  3.5-3.6
One Dimensional Heat Equation 8  Fall Break 9
3.7  Two Dimensional Equations, Midterm 2 10 3.8-3.9
Laplaces and Poissons Equations 11 4.1-4.2 Circular Coordinate and Vibration 12
4.3-4.4  Laplaces Equation in Circular Regions 13 4.4
Laplaces Equation in Circular Regions, Midterm 3 14 7.1  Fourier Integral Representation 15
7.2-7.3 Fourier Transform 16 7.4-7.5

Heat Kernel and Poisson Integral Formula